OpenStudy (anonymous):

The dimension of the edges in a right rectangular prism have the ratio 3:7:13. The volume of the prism is 1350 cm3. What is the length of the shortest side? (Round your answer to one decimal place.)

OpenStudy (anonymous):

A ratio of 3:7:13 can be represented algebraically as 3x:7x:13x. The volume of a prism is length x width x height Since the volume is 1350, your equation is (3x)(5x)(13x) = 1350 Try solving this equation.

OpenStudy (anonymous):

6.9

OpenStudy (anonymous):

x^3 = 6.92308 so what is x equal to?

OpenStudy (anonymous):

2.3076

OpenStudy (anonymous):

No. I get x = 1.9059

OpenStudy (anonymous):



OpenStudy (anonymous):

So the length of the shortest side is what? The shortest side will be the side represented by 3x. So what's the length of the shortest side?

OpenStudy (anonymous):

@Pash007 ?

OpenStudy (anonymous):

3(1.9059) = 5.7177

OpenStudy (anonymous):

Corrected. So 5.7 to the nearest tenth (as you directions asked to round to one decimal place).

OpenStudy (anonymous):

i have a question... In the first part of the equation.. (3x)(5x)(13x) = 1350 , why wasn't 7x used?

OpenStudy (anonymous):

Oh my...I read that 7 as a 5...so (3x)(7x)(13x) = 1350 sorry, my error in reading the numbers correctly.

OpenStudy (anonymous):

so 273x^3 = 1350 x^ 3 = 4.94505 x = 1.703 so shortest side is 3(1.7030) = 5.1 to nearest tenth. My apologies. @Pash007